# Quantum Machine Learning and Grover’s Algorithm for Quantum Optimization of Robotic Manipulators

### v1

@@ V-A Implementation Framework
The quantum optimization algorithm was implemented using Qiskit [34] and executed on IBM’s ibm_brisbane superconducting quantum processor (with circuit depths up to 100 gates and 36 qubits in tested cases). The framework integrates QML with Grover’s algorithm, where the QML circuit learns the forward kinematics model and dynamically constructs the oracle based on the cost function, while Grover’s algorithm amplifies states associated with optimal configurations. Circuits were executed on real hardware after initial validation on the ibmq_qasm_simulator, with transpilation respecting topology and coherence limits. Although IonQ hardware was not used here, the design is compatible with trapped-ion processors, and future work will evaluate scalability on such platforms. Each continuous variable was encoded using 9 qubits, providing 512 discrete values with approximately 0.7° angular resolution. Circuit transpilation respected hardware topology constraints, with error mitigation techniques applied to enhance results. The complete optimization procedure is summarized in Algorithm 2.
Data: Target pose $\bm{p}_{\text{target}}$, discretized state space $\{\lvert k\rangle\}$
Result: Optimal configuration $\lvert k^{*}\rangle$
Prepare initial superposition state $\lvert\psi\rangle=\sum_{k}c_{k}\lvert k\rangle$;
Train QML circuit $U_{\text{QML}}$ to compute forward kinematics $f(\cdot)$;
Construct oracle $O$ to mark states where $C(k)=\left\|f(k)-\bm{p}_{\text{target}}\right\|^{2}<\epsilon$;
for $k=1$ to $K$ do
   Apply oracle $O$ and diffusion operator $D$;
Measure final state to retrieve optimal $\lvert k^{*}\rangle$;
return Decoded configuration from $\lvert k^{*}\rangle$;
Algorithm 2 Quantum Optimization for Inverse Kinematics
@@ V-B One DoF Case
Figure 3: Quantum variational circuit for optimizing a single-link manipulator to generate a circular workspace. The circuit uses parameterized rotations ($R_{X}$, $R_{Y}$) and CNOT gates across 9 qubits to encode solutions and evaluate their fitness through measurement, enabling iterative refinement of the joint angle. It is tailored to optimize the cost function, incorporating link length and other relevant parameters.
The quantum search determined the optimal link length $l_{1}^{*}=1$, at which the end-effector’s workspace coincides with the target manifold ($S^{1}$) within the error threshold. For this length, the framework solved the IK for target point $\bm{p}=[0;1]$. The algorithm amplified the state corresponding to $\theta_{1}=90^{\circ}$, aligning the manipulator vertically along the positive y-axis. The quantum approach identified optimal configurations, but total time (2.0s) exceeded classical optimization (1.1s for Nelder-Mead) for this case, resulting in a 0.55x speedup (i.e., quantum is slower) as in Fig. 7. This reflects quantum overhead for low-dimensional problems where classical methods are efficient. However, it demonstrates the method’s capability, establishing a baseline for complex implementations.
@@ V-C Two DoF Case
The framework was applied to a 2-DoF planar manipulator to determine optimal link lengths $l_{1}$, $l_{2}$ and joint angles $\theta_{1}$, $\theta_{2}$ enabling the end-effector to reach the target position within the configuration space topology illustrated in Fig. 2. Each parameter was discretized into 512 levels using 9 qubits, resulting in a search space of $512^{4}$ configurations and 36 qubits. The comparative analysis (Fig. 7) demonstrates quantum advantage. Optimization converged in 22.5 seconds with 68% success probability after mitigation. Classical methods took longer: BFGS 2600 seconds, Nelder-Mead 2100 seconds, PSO 2200 seconds—a 93.3x speedup over Nelder-Mead. The approach identified optimal $l_{1}$ and $l_{2}$ reaching the target within the toroidal configuration space with constraints. Fig. 5(a) shows cost convergence, and Fig. 5(b) shows link length progression. This advantage stems from Grover’s speedup in discrete spaces, where classical exhaustive evaluation requires extensive time. The results highlight efficiency for multi-parameter optimization, with advantages in higher dimensions.
Figure 4: Optimization results for one- and two-DoF manipulators. (a) Quantum-based cost function convergence for the one-DoF case. (b) Quantum-based joint-angle trajectory for the one-DoF case. (c) Classical cost convergence for the two-DoF case. (d) Classical link-length convergence for the two-DoF case.
Figure 5: Optimization results for the 2-DoF manipulator (quantum) and the dual-arm system (classical). (a) Quantum cost convergence for the 2-DoF case. (b) Quantum link-length updates. (c) Schematic of dual-arm grasping with contact points on a 2D sphere. (d) Classical grasping cost convergence based on contact-point error.
Figure 6: Optimization for dual-arm. (a) Classical joint-angle trajectories for grasping. (b) Quantum convergence of the contact-point distance cost. (c) Quantum evolution of optimal joint angles.
@@ V-D Dual-Arm Grasping
The quantum implementation, following the five-stage pipeline outlined in Section IV-C, employed a total of $N=36$ qubits to encode the discretized joint space. The key result is a demonstrated 9.4× speedup over classical exhaustive search methods in identifying optimal configurations, attributed to the quadratic acceleration provided by Grover’s algorithm. This performance gain was realized without compromising solution quality.
Classical verification of the quantum-derived solutions confirmed that the grasping error satisfied $e_{\text{actual}}\leq\epsilon$. All decoded joint configurations $(\bm{q}_{1}^{*},\bm{q}_{2}^{*})$ were also validated to be feasible, satisfying collision avoidance and force-closure conditions. This case study thus confirms the framework’s capacity and scalability for complex, multi-robot coordination tasks, marking a significant step toward quantum-accelerated solutions in collaborative robotics.
@@ V-E Comparative Performance Analysis
Figure 7: Performance comparison across manipulator configurations and optimization methods.
The computational performance across different manipulator configurations, summarized in Fig. 7, demonstrates the quantum framework’s distinctive scaling properties. For the 1-DoF case, quantum optimization required 2.0 seconds compared to 1.1 seconds for the classical Nelder-Mead approach, resulting in a 0.55x speedup (indicating quantum overhead dominates for simple problems). This overhead includes state preparation, oracle construction, and measurement operations that become negligible as problem dimensionality increases. Substantial quantum advantage emerges in higher-dimensional configurations: the 2-DoF manipulator optimization achieved a 93.3x speedup (22.5 seconds quantum vs. 2100 seconds classical), while the dual-arm system demonstrated a 9.4x improvement. This performance scaling aligns with Grover’s theoretical $\mathcal{O}(\sqrt{N})$ complexity advantage over classical methods’ exponential scaling. The results show that quantum optimization surpasses classical methods as complexity increases, offering clear advantages for high-dimensional robotic applications.

### v2

@@ V-A Implementation Framework
The quantum optimization algorithm was implemented using Qiskit [32] and executed on IBM’s ibm_brisbane superconducting quantum processor (with circuit depths up to 100 gates and 36 qubits in tested cases). The framework integrates QML with Grover’s algorithm, where the QML circuit learns the forward kinematics model and dynamically constructs the oracle based on the cost function, while Grover’s algorithm amplifies states associated with optimal configurations. Circuits were executed on real hardware after initial validation on the ibmq_qasm_simulator, with transpilation respecting topology and coherence limits. Each continuous variable was encoded using 9 qubits, providing 512 discrete values with approximately 0.7° angular resolution. Circuit transpilation respected hardware topology constraints, with error mitigation techniques applied to enhance results.
@@ V-B One DoF Case
Figure 3: Quantum variational circuit (ansatz design) for optimizing a single-link manipulator to generate a circular workspace. The circuit uses parameterized rotations ($R_{X}$, $R_{Y}$) and CNOT gates across 9 qubits to encode solutions and evaluate their fitness through measurement, enabling iterative refinement of the joint angle. It is tailored to optimize the cost function, incorporating link length and other relevant parameters.
The quantum framework successfully identified optimal configurations that satisfied the precision constraints (0.1–0.5 mm positional accuracy), demonstrating the method’s capability for basic robotic optimization and establishing a foundational baseline for more complex implementations.
@@ V-C Two DoF Case
The framework was applied to a 2-DoF planar manipulator to determine optimal link lengths $l_{1}$, $l_{2}$ and joint angles $\theta_{1}$, $\theta_{2}$ enabling the end-effector to reach the target position within the configuration space topology illustrated in Fig. 2. Each parameter was discretized into 512 levels using 9 qubits, resulting in a search space of $512^{4}$ configurations and 36 qubits.
The quantum optimization converged in 22.5 seconds with 68% success probability after mitigation. The approach successfully identified optimal $l_{1}$ and $l_{2}$ values that reach the target within the toroidal configuration space constraints. Fig. 5(a) shows the cost function convergence, and Fig. 5(b) illustrates the link length progression. The rapid convergence observed stems from Grover’s speedup in discrete spaces, where classical exhaustive evaluation becomes computationally prohibitive. These results demonstrate the method’s efficacy for multi-parameter optimization in robotic systems.
(a) Quantum-based cost function convergence for the one-DoF case over optimization iterations.
(b) Quantum-based joint-angle trajectory for the one-DoF case over optimization iterations.
(c) Classical cost convergence for the two-DoF case over optimization iterations.
(d) Classical link-length convergence for the two-DoF case over optimization iterations.
Figure 4: Optimization results for one- and two-DoF manipulators. (x-axes represent optimization iterations; figures regenerated with continuous lines without markers for clarity.)
(a) Quantum cost convergence for the 2-DoF case over optimization iterations.
(b) Quantum link-length updates for the 2-DoF case over optimization iterations.
(c) Schematic of dual-arm grasping with contact points on a 2D sphere.
(d) Classical grasping cost convergence based on contact-point error over optimization iterations.
Figure 5: Optimization results for the 2-DoF manipulator (quantum) and the dual-arm system (classical). (x-axes represent optimization iterations; figures regenerated with continuous lines without markers for clarity.)
(a) Classical joint-angle trajectories for grasping over optimization iterations.
(b) Quantum convergence of the contact-point distance cost over optimization iterations.
(c) Quantum evolution of optimal joint angles over optimization iterations.
Figure 6: Optimization for dual-arm. (x-axes represent optimization iterations; figures regenerated with continuous lines without markers for clarity.)
@@ V-D Dual-Arm Grasping
The quantum implementation, following the five-stage pipeline outlined in Section IV-C, employed a total of $N=36$ qubits to encode the discretized joint space. Figs. 5(d) and 6(a) show the classical optimization results for the dual-arm system shown in Fig. 5(c), depicting the convergence of the grasping cost based on contact-point error and the trajectories of joint angles over optimization iterations, respectively, while Figs. 6(b) and 6(c) illustrate the quantum-based results. The key result is a demonstrated 9.4× speedup over classical exhaustive search methods in identifying optimal configurations, attributed to the quadratic acceleration provided by Grover’s algorithm. This performance gain was realized without compromising solution quality, as illustrated in the convergence plots depicted in Fig. 5(c).
@@ V-E Comparative Performance Analysis
We benchmark against three classical optimizers representing distinct optimization approaches: BFGS [33] (gradient-based), Nelder-Mead [34] (direct search), and PSO [35] (population-based metaheuristic). This selection enables robust cross-framework comparison across different problem characteristics and complexity levels. Figure 7 summarizes computational performance across manipulator configurations. For the 1-DoF case, the quantum approach required 2.0 seconds compared to 1.1 seconds for Nelder-Mead (0.55x speedup), indicating quantum overhead dominates for simple problems where classical methods are inherently efficient. However, this overhead becomes negligible as dimensionality increases.
For the 2-DoF case, quantum optimization achieved 93.3x speedup (22.5 seconds vs. 2100 seconds for Nelder-Mead). Other classical methods required 2600 seconds (BFGS) and 2200 seconds (PSO). The dual-arm system demonstrated a 9.4x improvement, confirming the scaling trend.
Figure 7: Performance comparison showing quantum advantage scaling with problem dimensionality.
This performance pattern aligns with Grover’s $\mathcal{O}(\sqrt{N})$ complexity advantage over classical exponential scaling. The quantum overhead includes state preparation, oracle construction, and measurement operations that become amortized in higher-dimensional problems.
Experimental parameters were carefully designed to navigate the trade-offs between current NISQ device limitations and practical industrial robotic requirements. The 9-qubit discretization scheme per parameter was selected to provide sufficient resolution for high-precision applications while remaining executable on available quantum hardware. This encoding strategy enables sub-millimeter positional accuracy (0.1–0.5 mm) and sub-degree angular precision (0.1–0.7°), meeting stringent industrial standards for applications such as precision assembly, micro-manufacturing, and surgical robotics. The circuit depth and gate complexity were optimized to accommodate current coherence time constraints while maintaining computational accuracy. Error mitigation techniques were employed to address quantum decoherence and gate infidelities inherent in contemporary quantum processors. The consistent performance scaling observed across problem dimensionalities—from simple 1-DoF cases to complex multi-arm systems—validates the practical viability of this approach for real-world precision engineering applications, demonstrating that quantum optimization can deliver both theoretical advantages and measurable improvements in high-accuracy robotic tasks.